OpenStudy (anonymous):

Hi all! I need help visualizing this question. :) "Express the surface area of a cube as a function of its volume"

OpenStudy (campbell_st):

well whats the volume if the side length is x whats the surface area of the same cube, with side length x thats the start of the question...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

V=x^3 ?

OpenStudy (anonymous):

Surface area= 6x^2 ?

OpenStudy (anonymous):

Is this right?

OpenStudy (nurali):

Welcome to Openstudy @grinnell.12

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

I'm a bit confused still I must say

OpenStudy (anonymous):

I'm trying to see how I should think of this in general

OpenStudy (anonymous):

"express s in terms of V." what does that really mean?

OpenStudy (anonymous):

"so s = V^(1/3)" this is the part I am not understanding :(

OpenStudy (anonymous):

why do you take the cube root?

OpenStudy (anonymous):

Is it that you are taking V = s^3 and solving for s?

OpenStudy (anonymous):

so that's the "express s in terms of V."

OpenStudy (anonymous):

in terms of just means using the equation for V

OpenStudy (anonymous):

?

OpenStudy (anonymous):

great thanks again :)